Circular No. 7247

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999dp IN UGC 3046 W. D. Li, University of California at Berkeley, on behalf of the Lick

Observatory Supernova Search (LOSS; cf. IAUC 6627, 7126), reports the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT), of an apparent supernova in UGC 3046. SN 1999dp was discovered on an unfiltered image taken on Sept. 2.5 UT and was confirmed on earlier images taken on Aug. 18.5, 25.5, and 28.5, all showing the star at mag ∼ 18.2. The new object is located at α = 4^{h}29^{m}16^{s}.66, δ = +69^{o}32′00′′.6 (equinox 2000.0), which is 18′′.0 west and 7′′.8 north of the nucleus of UGC 3046. There is a nearby star (mag ∼ 16.5) with position end figures 15^{s}.87, 02′′.4. A KAIT image of the same field on 1998 Aug. 31.5 (limiting magnitude ∼ 19.5) showed nothing at the position of SN 1999dp.

SUPERNOVA 1999dq IN NGC 976 Li also reports the discovery by the LOSS of an apparent supernova in

NGC 976. SN 1999dq was discovered and confirmed on unfiltered images taken on Sept. 2.5 (mag ∼ 16.3) and 3.4 UT (mag ∼ 16.0). The new object is located at α = 2^{h}33^{m}59^{s}.68, δ = +20^{o}58′30′′.4 (equinox 2000.0), which is 4′′.3 west and 6′′.4 south of the nucleus of NGC 976. KAIT images of the same field on Aug. 27.5 (limiting mag ∼ 17.0) and 24.5 (limiting mag ∼ 19.0) showed nothing at the position of SN 1999dq.

SAX J1712.6–3739 M. Cocchi and L. Natalucci, Istituto di Astrofisica Spaziale, CNR,

Rome; J. in ’t Zand and J. Heise, Space Research Organization Netherlands (SRON), Utrecht; J. M. Muller, BeppoSAX, Rome, and SRON; and G. Celidonio and L. Di Ciolo, BeppoSAX, Rome, communicate: “During an observation with Wide Field Camera 2 onboard BeppoSAX on Sept. 1.68– 3.16 UT, an ∼ 20-s long x-ray burst was detected on Sept. 2.87 from a position 1′.9 away from SAX J1712.6–3739 (IAUC 7243). The 99-percent-confidence error radius is 3′.5. The peak flux on a time scale of 1 s is 1.7 Crab units (2–26 keV). This burst and its time profile (fast rise and exponential decay) suggest that SAX J1712.6–3739 is a low-mass x-ray binary with a neutron star as compact object. No steady emission was detected above an upper limit of 7 mCrab over the whole observation. During an observation with the narrow-field instruments of BeppoSAX on Aug. 27.43–27.95, the flux was 6 mCrabs (2–10 keV).”

1999 September 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green